# Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation

### v1

@@ Appendix G Baseline Descriptions
  • ChatGLM Zeng et al. (2022) is an open-source and bilingual language model, based on GLM Du et al. (2021) architecture. The technique of ChatGLM-6B is similar to ChatGPT, optimized in dialogue tasks.
  • Flan-T5 Chung et al. (2022) proposed several instruction-based LLMs, scaling the number of tasks and model size and fine-tuning in the chain-of-thought data. We adopt Flan-T5-XXl (11B), Flan-T5-Large (780M) and Flan-T5-Small (80M) as baselines in our work.
@@
  • BM25 Robertson et al. (2009): It is the ranking function between a query and several documents widely used in information retrieval. In our experiments, we also apply RefinED to retrieve triples for BM25 and replace documents with triples for retrieval.
  • KAPING Baek et al. (2023): It is a popular knowledge-augmented method for question answering. Since the authors have not released the official code, we follow the experimental setup of the KAPING method: we first extract entities from the query using REfinED, rank the query with triples using MPNet, select the top-N most relevant triples, and supplement them as prompts to generate dialogue responses.
  • G-Retriever He et al. (2024): It is a retrieval-augmented generation method for textual graphs that selects informative subgraphs via a Prize-Collecting Steiner Tree, encodes them with GNNs, and generates responses with high accuracy, outperforming strong baselines and reducing hallucinations.
  • Triton(Ours): We use GPT-4o to generate high-quality query-triple pairs based on the training dataset and fine-tune Llama3 8B base model with collected samples. The model acts as a triple retriever to help select valuable triples for generating dialogue responses.
  • Triton-C(Ours): We fine-tune a coreference resolution model based on Llama 8B base model using reasoning distillation. Our approach integrates dialogue rewriting with entity linking for enhanced performance.
  • Triton-X(Ours): It is the framework that consists of our well-designed knowledge retriever, dialogue rewrite and graph-based module.
Appendix H Details of the Case Study
The examples of case studies are shown in Table 12.
The topic in Case 1 is about the population of Montevideo, and our methods do help select the relevant triples. However, the LLM then generates one of triples directly, which is not fluent and has a higher PPL i.e. lower quality.
Case 2 is a dialogue that asks about the release date of “Team Fortress 2”, and the original response responds to the dialogue and is close to the writing of reference. After rewriting the dialogue, the LLM generates a response that differs from the reference, causing a lower BLEU-4 score.

### v2

@@ Appendix F Baseline Descriptions
  • ChatGLM (Zeng et al. 2022) is an open-source and bilingual language model, based on GLM (Du et al. 2021) architecture. The technique of ChatGLM-6B is similar to ChatGPT, optimized in dialogue tasks.
  • Flan-T5 (Chung et al. 2022) proposed several instruction-based LLMs, scaling the number of tasks and model size and fine-tuning in the chain-of-thought data. We adopt Flan-T5-XXl (11B), Flan-T5-Large (780M) and Flan-T5-Small (80M) as baselines in our work.
@@
  • BM25 (Robertson, Zaragoza et al. 2009): It is the ranking function between a query and several documents widely used in information retrieval. In our experiments, we also apply RefinED to retrieve triples for BM25 and replace documents with triples for retrieval.
  • KAPING (Baek, Aji, and Saffari 2023): It is a popular knowledge-augmented method for question answering. Since the authors have not released the official code, we follow the experimental setup of the KAPING method: we first extract entities from the query using REfinED, rank the query with triples using MPNet, select the top-N most relevant triples, and supplement them as prompts to generate dialogue responses.
  • G-Retriever (He et al. 2024): It is a retrieval-augmented generation method for textual graphs that selects informative subgraphs via a Prize-Collecting Steiner Tree, encodes them with graph neural networks, and generates responses with high accuracy, outperforming strong baselines and reducing hallucinations.
Appendix G Dialogue Reformulation Evaluation
We evaluate the dialogue reformulation performance through human judgment, with annotation details provided in  E. Each instance is rated as good (1), neutral (0.5), or bad (0), and the final score is computed as the average. Our method achieves a score of 0.9325 on the HybriDialogue dataset and 0.9475 on the OpenDialKG dataset, demonstrating the strong effectiveness of our proposed reasoning-guided dialogue reformulation approach.
Appendix H Addtional Case Studies
The additional case studies are shown in Table 13.
Case 2 presents a dialogue about the book Blinky Bill. The original dialogue contains coreference issues, making the dialogue difficult to follow and resulting in a factually incorrect response. By applying our TGDRG-R approach, which combines step-by-step reference resolution in reasoning-guided dialogue reformulation with dialogue sense knowledge selection, we achieve the generation of correct and coherent responses.
The topic in Case 3 is about the population of Montevideo, and our methods do help select the relevant triples. However, the TG-DRG-NR then generates one of the triples directly, which is not fluent and has a higher PPL i.e. lower quality.